13. Selling price and the amount spent on advertising were entered into a multiple regression to determine what affects flat panel LCD TV sales. Using the output below, which statement(s) is(are) correct? Analysis of Variance Source DF (Adj)SS MS Regression 2 16477.3 8238.7 Error 27 3038.0 112.5 Total 29 19515.4 I. The F statistic to determine the overall significance of the estimated multiple regression model is 73.23. II. The multiple regression model is overall significant. III. It is not appropriate to use multiple regression model. A) I only B) III only C) Both I and II D) Both I and III E) II only
Answer #1

here F test statisitc =MSR/MSE =8238.7/112.5 =73.23

as F test statistic is in critical region the model is significantt

option C is correct

C) Both I and II